# ImpMIA: Leveraging Implicit Bias for Membership Inference Attack under Realistic Scenarios

### v1

@@ 6.2 Assumptions Influence Analysis
We present a detailed analysis of the influence of removing each assumption—individually and jointly—on attack performance. As discussed earlier, reference-model–based methods rely on training reference models under conditions that closely match the target model. Their effectiveness comes from approximating the loss distributions of members versus non-members, which requires alignment between target and reference training.
@@
In Table 2 we report results on the CINIC-10 dataset across five scenarios: all assumptions provided (”Standard Setting”), the removal of each assumption individually, and the removal of all three assumptions simultaneously. As shown, all methods except our proposed ImpMIA—which does not rely on any reference models—suffer substantial drops in performance when all assumptions are removed. In particular, LiRA and Attack-R exhibit a clear degradation as assumptions are removed. While they perform well in the standard setting, their TPR at low FPR decrease substantially under unknown configuration, distribution shift, and unknown ratio. In contrast, our method achieves strong results in these settings, as it does not depend on such assumptions. We observed improvements in ImpMIA at $0.01\%$ FPR level when the fraction of members was changed, which may be explained by the evaluation metric’s dependence on the pool size (see Appendix B.6). It is worth noting that RMIA’s performance in the standard setting is lower than those reported in its original paper. This is because our training follows the white-box setup from Cohen & Giryes (2024), which differs from RMIA’s original experimental setting. This may imply limited generalization of RMIA (see Appendix C.1 for further discussion).
Figure 3 further visualizes the effect of assumption removal (for 0% FPR) in CINIC-10 dataset, for the top-performing method (LiRA) and our ImpMIA. As seen, being a reference-model–based method, LiRA degrades sharply as assumptions are removed, while ImpMIA shows no loss in performance. Overall, these results demonstrate the robustness and effectiveness of our approach for membership inference under realistic conditions where the attacker cannot rely on privileged training knowledge.
7 Discussion & Limitations
ImpMIA advances membership inference by introducing a simple, theory-driven white-box attack that outperforms both black- and white-box baselines under realistic scenarios. By avoiding the need for reference models and leveraging implicit bias in neural network training, our approach provides a more practical and scalable way to audit privacy in machine learning systems. The main assumptions of our attack are access to model weights and to a candidate superset containing the training set. The first assumption is increasingly realistic as many modern models are publicly released with their full parameters, and the second is reasonable since our method’s efficiency allows scaling to very large candidate pools without the need to train reference models (see Section B.2). While performance is strongest when most of the training set is included, we show that our attack also works well with partial training coverage (see Section B.3).
@@ 7 Discussion & Limitations
Our work is the first to demonstrate the implications of implicit bias theory for membership inference attacks. More broadly, it provides a concrete case study of how insights from implicit bias theory, which have largely been developed in idealized or small-scale settings, can be instantiated in practical machine learning tasks, going beyond theory and toy examples to real datasets, larger neural networks, and standard training regimes. ImpMIA provides a step forward in practical privacy auditing, establishing a bridge between implicit bias theory and applied machine learning.
@@ Appendix A ImpMIA Additional Implementation Details
Finally, in the post-processing step, we refine the scores through margin-based boosting and distance scaling. Margin-based boosting increases scores for classes with lower average margins (harder classes), as these are more likely to contain misclassified test samples and hence stronger membership signals. Per-sample boosting also increases scores for points closer to the decision boundary, since such samples are more likely to be memorized. Distance scaling further penalizes deviation from the estimated class margin: we approximate the margin $\bar{m}_{c}$ as the mean margin of the top-$k$ highest-coefficient samples in each class, and rescale scores by dividing with $|\Delta_{i}-\bar{m}_{c}|^{\eta}$. This ensures that only samples whose margins align with expected member behavior maintain high scores, reducing false positives among non-members.
Training our proposed attack (ImpMIA) takes about 24 hours on a single H100 GPU, or 3 hours when distributed across 8 GPUs. In contrast, black-box reference-model attacks require about 8 days on a single GPU, or 1 day on 8 GPUs. The target models achieved average test accuracies of 90.5% on CIFAR-10, 74% on CINIC-10, and 66% on CIFAR-100.
Appendix B Ablation and Analysis
@@
In this section, we provide further ablations and analyses of our approach: (i) visualization of the $\lambda$ scores across different classes (see Section B.1); (ii) the ability to apply our attack on large candidate pools (see Section B.2) ; (iii) the effect of the number of training samples included in the superset (see Section B.3); (iv) ablations on the different scoring variants used in our method (see Section B.4); (v) the influence of weight decay on the results, showing that our method also works without explicit weight decay (see Section B.5); and (vi) the effect of candidate pool size on evaluation (see section B.6)
B.1 Visualization of $\lambda$ Scores
@@
In Figure S1, we present results for six different CIFAR-100 classes. The plots show the $\lambda$ score on the y-axis and the distance from the margin on the x-axis. As expected, high $\lambda$ scores are strong indicators of membership. Samples very close to the margin are more likely to be non-members, while almost all members with high $\lambda$ values fall slightly farther from the margin. This is consistent with the fact that models are trained to push training samples away from the margin, while hard test samples can lie closer to it. Interestingly, training samples that remain close to the margin are those that the network tends to memorize (Haim et al., 2022), and therefore receive higher $\lambda$ scores, reflecting their strong influence on the model’s predictions and weights.
@@
Figure S1: Lambda Scores Visualization (CIFAR-100). Scatter plots for six different classes. Each plot shows superset samples; x-axis is distance from the decision boundary, y-axis is the $\lambda$ score, and points are colored by membership (member vs. non-member). High $\lambda$ values strongly indicate membership.
@@ B.2 Large Candidate Pool
We further evaluated ImpMIA under settings with larger candidate pools. In our previous experiments, we used candidate pools of 50,000 ($\approx 2\times$ the training size) and 80,000 ($\approx 3\times$ the training size). To stress scalability, we also ran experiments with a candidate pool of 250,000 samples ($\approx 10\times$ the training size). Specifically, we trained the target model on 25,000 samples from the CINIC-10 dataset (combining all three splits) and used 250,000 candidate samples drawn from the remainder of the dataset ($\approx 270{,}000$ images in total). To make these large-scale experiments feasible, we introduced two adaptations: (i) reduced precision, where gradient batches were computed in FP16 instead of FP32 to lower memory requirements; and (ii) filtered optimization, where candidates were filtered by estimating each class’s distance-to-boundary distribution, identifying the peak of this distribution, and retaining only the samples within a small interval around the peak. Despite the much larger candidate pool, our results remained strong, achieving a TPR of $5.49\%$ at $0.01\%$ FPR, $0.97\%$ at $0.00\%$ FPR, and an AUC of $0.79$.
B.3 Effect of Training Sample Coverage in the Superset
Our method relies on the implicit bias of the network, linking training samples to the learned weights. Consequently, if a large portion of the training set is missing from the candidate superset, performance naturally decreases. Importantly, in practice it is reasonable to expect supersets that cover most of the training set, especially since our method’s efficiency allows scaling to very large candidate pools without the need to train reference models. As shown in Table T1, performance is strongest when most of the training set is included, but our attack also works well under partial training coverage. In all cases, evaluation is on a random $5,000$ ($10\%$) samples from the candidate pool to avoid influence from candidate pool size on the reported metrics.
Table T1: Effect of Training Sample Coverage. Ablation study on CIFAR-10 showing the impact of training sample coverage within the candidate superset. Performance is strongest when most of the training set is included, but our method remains effective under partial coverage.
@@ B.3 Effect of Training Sample Coverage in the Superset
100% 0.72 5.13 2.63
75% 0.71 3.63 2.20
50% 0.70 3.10 1.45
25% 0.66 2.15 1.21
10% 0.62 1.54 0.39
B.4 Scoring Variants Ablation
We present an ablation study of the different score-refinement components in our pipeline (see Table T2). After block-wise optimization, each sample has multiple coefficient estimates—one per block—which are fused into a robust score using both a trimmed mean (to discard extreme outliers) and a signal-to-noise ratio (SNR, mean over standard deviation across blocks). This aggregation step already improves robustness by emphasizing consistent membership signals. On top of this, we evaluate margin-based boosting and distance scaling. Class-level boosting gives higher weight to harder classes (with lower average margins), while per-sample boosting highlights points near the decision boundary, which are more likely to be memorized. Finally, distance scaling penalizes samples whose margins deviate from the estimated class margin, reducing false positives. As shown, each component contributes to performance gains, and the full ImpMIA pipeline achieves the strongest low-FPR detection.
Table T2: Effect of Score Aggregation and Post-Processing. Ablation study on CIFAR-100 analyzing coefficient aggregation, margin-based boosting, and distance scaling. After optimizing each block separately, coefficients are aggregated across blocks using a trimmed mean (central values only) and a signal-to-noise ratio (SNR). Margin boosting increases scores for harder classes (lower margins) and samples near the decision boundary, while distance scaling penalizes deviations from the estimated class margin. Together, these refinements yield the final ImpMIA scores.
@@ B.4 Scoring Variants Ablation
Trimmed mean only 0.86 8.20 5.82
Robust SNR only 0.90 4.63 3.60
Fusion (trimmed mean + SNR) 0.91 8.24 5.86
+ Boost1 (class-level margins) 0.91 8.33 5.91
+ Boost2 (sample-level margins) 0.90 9.27 6.55
+ Trim division 0.91 8.85 6.36
ImpMIA 0.90 9.66 6.73
B.5 Weight Decay Influence
To study the effect of explicit weight decay, we evaluated ImpMIA across several decay levels as well as without decay. While the KKT stationarity formulation applies in the homogeneous case, prior theory suggests that in non-homogeneous networks explicit weight decay is needed as a replacement. However, our results show that even in this setting the attack remains effective without weight decay. Interestingly, smaller decay values (e.g., $10^{-6}$, $10^{-5}$) tend to improve performance compared to larger ones, reflecting a balance between stability and memorization. Moreover, the case without weight decay performs comparably—and in some metrics slightly better—than with decay, consistent with the intuition that stronger regularization reduces memorization and weakens membership signals. This demonstrates both the robustness of our method and that weight decay is not strictly necessary for strong empirical performance.
Table T3: Effect of Weight Decay. Ablation on CIFAR-10 showing performance across different levels of weight decay, as well as without it. Although implicit bias theory assumes weight decay for the KKT characterization, our results show the attack remains effective regardless. Smaller decay values yield stronger performance, and even the no–weight decay case performs competitively, likely because stronger decay suppresses memorization signals.

### v3

@@ 6.2 Assumptions Influence Analysis
We present a detailed analysis of the influence of removing each assumption, individually and jointly, on attack performance. As discussed earlier, reference-model-based methods rely on training reference models under conditions that closely match the target model. Their effectiveness comes from approximating the loss distributions of members versus non-members, which requires alignment between target and reference training.
@@
In Tab. 3 we report results on the CINIC-10 dataset across five scenarios: all assumptions provided (Full Auxiliary-Knowledge setting), the removal of each assumption individually, and the removal of all three assumptions simultaneously. As shown, all methods except our proposed ImpMIA, which does not rely on any reference models, suffer substantial drops in performance when all assumptions are removed. In particular, LiRA and Attack-R exhibit a clear degradation as assumptions are removed. While they perform well in the Full Auxiliary-Knowledge setting, their TPR at low FPR decrease substantially under unknown configuration, distribution shift, and unknown ratio. In contrast, our method achieves strong results in these settings, as it does not depend on such assumptions. We observed improvements in ImpMIA at $0.01\%$ FPR level when the fraction of members was changed, which may be explained by the evaluation metric’s dependence on the pool size (see  Sec. B.4). It is worth noting that RMIA’s performance in the Full Auxiliary-Knowledge setting is lower than those reported in its original paper. This is because our training follows the white-box setup from Cohen and Giryes (2024), which differs from RMIA’s original experimental setting. This may imply limited generalization of RMIA (see  Sec. C.1 for further discussion).
6.3 Ablation and Analysis
The Appendix presents additional ablation studies and analyses. This includes: visualizations of the $\lambda$ values (Sec. B.1); applicability to large candidate pools (Sec. B.2); performance under partial training-set coverage (Sec. B.3); the effect of pool size on evaluation (Sec. B.4); ablations on the scoring and aggregation components (Sec. B.5), the effect of weight decay (Sec. B.6) and pre-filtering strategies (Sec. B.7), and running time analysis (Sec. B.8).
7 Conclusion
@@
ImpMIA advances Membership Inference by introducing a simple theory-driven white-box attack which leverages the implicit bias in neural-nets. It outperforms both black- and white-box baselines in scenarios where the attacked network’s training procedure & settings are unknown. By resorting to pure optimization of KKT conditions on the attacked model’s parameters (without the need to train any reference models), our method allows efficient scaling to very large candidate pools (see Sec. B.2). The white-box assumption is increasingly realistic as many modern models are publicly released with their full parameters. ImpMIA is best suited for scenarios where the training-set is contained in the examined-set. However, we show that this assumption can be relaxed (Sec. B.3), and that ImpMIA is competitive even at low coverage of the training-set (only 10%), and becomes superior at high coverage. Our work is the first to demonstrate the implications of implicit bias theory to membership inference attacks. More broadly, it provides a concrete case study of how insights from implicit bias theory, which have largely been developed in idealized or small-scale settings, can be instantiated in practical machine learning tasks, going beyond theory and toy examples to real datasets, larger neural networks, and standard training regimes. ImpMIA provides a step forward in practical privacy auditing, establishing a bridge between implicit bias theory and applied machine learning.
Acknowledgments
This research was funded by the European Union (ERC grant No. 101142115). GV is supported by the Israel Science Foundation (grant No. 2574/25), by a research grant from Mortimer Zuckerman (the Zuckerman STEM Leadership Program), and by research grants from the Center for New Scientists at the Weizmann Institute of Science, and the Shimon and Golde Picker – Weizmann Annual Grant.
@@ Appendix A ImpMIA Additional Implementation Details
Finally, in the post-processing step, we refine the raw scores using three components: (i) class-level margin boosting, (ii) sample-level margin boosting, and (iii) distance-based scaling. Both boosting mechanisms adjust a sample’s score according to its margin (distance from the decision boundary), but at different granularities (class vs. individual sample). (i) Class-level boosting is a simple per-class rescaling step, motivated by prior work (Nasr et al., 2019) showing that some classes are systematically more vulnerable to membership inference. Our boost builds on the observation that harder classes (with lower average margins) are more likely to memorize additional samples, so an attack can potentially recover more members from these classes. Accordingly, we assign slightly higher weight to harder classes and mildly downweight easier ones via a small per-class scaling coefficient of $0.03$. (ii) Sample-level margin boosting then refines scores within each class: following previous observations (e.g., Haim et al.) that near-boundary points are more likely to be memorized, we slightly upweight low-margin, near-boundary samples and mildly downweight very high-margin ones, using a per-sample scaling factor of $0.3$. (iii) Distance-based scaling further penalizes deviation from the estimated class margin. We approximate the class margin $\bar{m}_{c}$ as the mean margin of the top-$k$ highest-coefficient samples in each class, and rescale scores by dividing by $|\Delta_{i}-\bar{m}_{c}|^{\eta}$, where $\Delta_{i}$ is the margin of sample $i$ and $\eta>0$ controls how strongly this distance affects the rescaling (we set $\eta=0.05$ in all experiments). This ensures that only samples whose margins are consistent with typical member behavior retain high scores, helping to reduce false positives among non-members.
Layer Selection for CIFAR-100.
For CIFAR-10 and CINIC-10 we construct the gradient matrix $A$ from all convolutional layers of the network. For CIFAR-100, however, we restrict $A$ to the final convolutional stages. This choice is motivated by both prior work and empirical observations. Nasr et al. (2019) empirically observed that, in deep CNNs trained on CIFAR-100, the strongest membership signals reside in the gradients of the later layers, while earlier layers mostly learn generic low- and mid-level features that generalize well and leak relatively little membership information. In a many-class regime such as CIFAR-100, early and mid-level layers must be heavily shared across classes, whereas the final convolutional blocks and classifier head implement sharp, class-specific decision boundaries and absorb most of the memorization.
Hyperparameter Selection.
All hyperparameters of ImpMIA were selected using a held-out evaluation set. For CIFAR-10 and CINIC-10 we used a dedicated validation split constructed from CINIC-10, which is large enough to provide an independent validation pool. For CIFAR-100 we used a validation split taken from the generated candidate set. For each dataset, we followed the same protocol as in our main experiments: we trained 5 target models with different random seeds (affecting both the data split and the initialization) under the realistic No-Auxiliary-Knowledge setting, and ran ImpMIA for a small grid of hyperparameter configurations. We then selected the configuration that achieved the best average performance across these 5 runs according to our main evaluation criteria (TPR at low FPR). The same hyperparameters are used across all scenarios and across all architectures (ResNet-18, VGG16, and ResNet50).
Appendix B Additional Ablation and Analysis
@@
In this section, we provide further ablations and analyses of our approach: (i) visualization of the $\lambda$ scores across different classes (Sec. B.1); (ii) scalability to large candidate pools (Sec. B.2); (iii) the effect of training-set coverage in the evaluated set (Sec. B.3); (iv) the effect of the non-member ratio and pool size on evaluation (Sec. B.4); (v) ablations of the scoring and aggregation variants (Sec. B.5); (vi) the influence of weight decay (Sec. B.6); (vii) the influence of pre-filtering on performance (Sec. B.7); and (viii) an analysis of the running time (Sec. B.8).
B.1 Visualization of $\lambda$ Scores.
@@
In Fig. S1, we present results for six different CIFAR-100 classes. The plots show the $\lambda$ score on the y-axis and the distance from the decision boundary on the x-axis. As expected, high $\lambda$ scores are strong indicators of membership. Samples very close to the decision boundary are more likely to be non-members, while almost all members with high $\lambda$ values fall slightly farther from the decision boundary. This is consistent with the fact that models are trained to push training samples away from the decision boundary creating a margin, while hard test samples can lie closer to it. Interestingly, training samples that remain close to the margin are those that the network tends to memorize (Haim et al., 2022), and therefore receive higher $\lambda$ scores, reflecting their strong influence on the model’s predictions and weights.
@@
Figure S1: Lambda scores visualization (CIFAR-100). Scatter plots for six different classes. Each plot shows the evaluation samples; x-axis is distance from the decision boundary, y-axis is the $\lambda$ score, and points are colored by membership (member vs. non-member). High $\lambda$ values strongly indicate membership.
@@ B.2 Large Candidate Pool
We further evaluated ImpMIA under larger candidate pools. In addition to the 50,000 ($\approx 2\times$) and 80,000 ($\approx 3\times$) candidate sets reported in the main text, we consider a pool of 250,000 candidates ($\approx 10\times$ the training size) on CIFAR-10. Following our realistic no auxiliary knowledge protocol, the target model is trained on 25,000 CIFAR-10 samples, while the reference set consists of the remaining 25,000 CIFAR-10 samples together with 200,000 images from CINIC-10. As shown in Table T1, ImpMIA achieves TPRs of $3.00\%$ at $0.01\%$ FPR and $0.54\%$ at $0.0\%$ FPR, substantially outperforming other MIAs at low FPRs in the larger candidate pool scenario.
Table T1: Membership inference with a 250K candidate pool (CIFAR-10). We compare ImpMIA to other MIA baselines under the realistic No-Auxiliary-Knowledge setting. We report AUC and TPR at fixed false-positive rates (FPR = 0.01% and 0.0%).
Attack AUC @0.01 % @0.0 %
Attack-R 0.75 0.01 0.0
Attack-P 0.76 0.12 0.0
LiRA 0.80 0.54 0.11
Ours 0.79 3.00 0.54
B.3 Effect of Training Sample Coverage in the Evaluation Set
Our method relies on the implicit bias of the network, linking training samples to the learned weights. Consequently, if a large portion of the training set is missing from the evaluation set, performance naturally decreases. Importantly, in practice it is reasonable to expect evaluation set that cover most of the training set, especially since our method’s efficiency allows scaling to very large candidate pools without the need to train reference models. As shown in Tab. T2, for the No-Auxiliary-Knowledge setting performance is strongest when most of the training set is included, but our attack also works well under partial training coverage. In all cases, evaluation is on a random $8,000$ ($10\%$) samples from the candidate pool to avoid influence from candidate pool size on the reported metrics.
In Tab. T3 we show a direct comparison with the previous best methods under only 10% coverage in our realistic scenario. Even with this low coverage, our results are comparable to the best competitors, and with higher coverage our method clearly surpasses them (Tab. 1). Higher coverage is a likely scenario in many practical settings, such as auditing whether a specific dataset was used to train a model, or when a dataset is known to contain the training set but the exact subset used is unknown..
Table T2: Effect of training sample coverage (no auxiliary knowledge). Ablation study on CIFAR-10 showing the impact of training sample coverage within the candidate superset. Performance is strongest when most of the training set is included, but our method remains effective under partial coverage.
@@ B.3 Effect of Training Sample Coverage in the Evaluation Set
100% 0.81 6.80 2.26
75% 0.81 5.49 1.42
50% 0.81 5.34 1.61
25% 0.79 3.49 0.96
10% 0.75 3.01 0.94
Table T3: Comparison under limited coverage of 10% (no auxiliary knowledge). We compare our method on CIFAR-10 with previous attacks when the candidate superset covers only 10% of the training data in our realistic No-Auxiliary-Knowledge setting. Even under this low coverage, our results are comparable to the best competitors.
Attack AUC @0.1 % @0.0 %
Attack-P 0.76 0.38 0.06
Attack-R 0.74 1.73 0.32
LiRA 0.80 2.98 1.20
RMIA 0.80 2.43 0.14
Ours 0.75 3.01 0.94
B.4 Influence of Non-Member Ratio on Evaluation
Our assumption-removal analysis also showed the influence of candidate pool size on model performance. In this setting, we lowered the member ratio by adding more non-members to the superset (increasing from 25K to 55K non-members). While our method does not rely on any specific member-to-non-member ratio, adding more non-members can affect results in two ways: (i) it increases the number of samples, which may make it harder to classify members, and (ii) it directly affects the evaluation metric, since FPR is defined relative to the total number of non-member samples. The overall performance reflects the interaction of these two factors. On the one hand, a larger pool introduces more distractors, potentially reducing accuracy. On the other hand, at fixed FPR thresholds (e.g., 0.01%), a larger pool allows more absolute mistakes. For instance, with 25K non-members, 0.01% FPR corresponds to only two false positives, whereas with 55K non-members, it allows up to five. Since coefficients are not uniformly distributed, permitting more mistakes can yield a nonlinear gain in TPR. In practice, we observed that at 0.01% FPR, our method performed better with 55K non-members than with 25K. This suggests that the positive effect of the evaluation metric outweighs the negative impact of additional distractors. Importantly, at 0.0% FPR, where the metric is unaffected by pool size, the results remained stable, as expected.
B.5 Scoring Variants Ablation
We present an ablation study of the different score-refinement components in our pipeline for CIFAR-10 in our No-Auxiliary-Knowledge setting (see Tab. T4). After block-wise optimization, each sample has multiple coefficient estimates, one per block, which are fused into a robust score using both a trimmed mean (to discard extreme outliers) and a signal-to-noise ratio (SNR, mean over standard deviation across blocks). This aggregation step already improves robustness by emphasizing consistent membership signals. On top of this, we evaluate margin-based boosting and distance scaling. Class-level boosting gives higher weight to harder classes (with lower average margins), while per-sample boosting highlights points near the decision boundary, which are more likely to be memorized. Finally, distance scaling penalizes samples whose margins deviate from the estimated class margin, reducing false positives. As shown, each component contributes to performance gains, and the full ImpMIA pipeline achieves the strongest low-FPR detection. Moreover, with either aggregation option, and even without any post-processing, our attack already outperforms previous attacks and shows strong results. The post-processing steps have only a comparatively small influence at low FPR (e.g., TPR@0.0% ranges only from 1.23 to 1.41), while aggregation has a larger effect. This is expected, since aggregation determines how we exploit the main signal produced by the KKT-based optimization. Overall, the primary source of our gains comes from the implicit-bias-based optimization itself rather than from the specific post-processing.
Table T4: Ablation of ImpMIA components for CIFAR-10 (no auxiliary knowledge). Performance of different scoring variants and incremental boosts. We report AUC and TPR at fixed false-positive rates (FPR = 0.01% and 0.0%).
@@ B.5 Scoring Variants Ablation
Trimmed mean only 0.74 2.58 1.18
Robust SNR only 0.85 2.51 0.89
Fusion (trimmed mean + SNR) 0.82 2.60 1.23
+ Boost1 (class-level margins) 0.82 2.59 1.23
+ Boost2 (sample-level margins) 0.81 2.69 1.39
+ Trim division 0.82 2.84 1.31
ImpMIA 0.81 2.76 1.41
B.6 Weight Decay Influence
To study the effect of explicit weight decay, we evaluated ImpMIA on CIFAR-10 No-Auxiliary-Knowledge setting, both with the standard decay level ($10^{-4}$) and without decay. While the KKT stationarity formulation applies in the homogeneous case, prior theory suggests that in non-homogeneous networks explicit weight decay is needed as a replacement. However, our results show that even in this setting the attack remains effective without weight decay: the no-weight decay variant achieves the same AUC (0.81) and slightly stronger performance in the low-FPR regime (2.93 vs. 2.76 at 0.01% FPR; 1.62 vs. 1.41 at 0.0% FPR). This is consistent with the intuition that stronger regularization can suppress memorization and weaken membership signals. Overall, this demonstrates the robustness of our method and that weight decay is not strictly necessary for strong empirical performance.
Table T5: Effect of Weight Decay (no auxiliary knowledge). Ablation on CIFAR-10 comparing the standard weight decay level ($10^{-4}$) to training without weight decay. Although implicit bias theory assumes weight decay in the non-homogeneous case for the KKT characterization, our results show the attack remains effective regardless.